OpenStudy (anonymous):

A coin is tossed 5 times. what is the probability of the coin lands heads more than once?

OpenStudy (anonymous):

a coin has 2 sides. you have 1 out of 2 chances to get heads. so far i know you would calculate this.. \[\frac{ 1 }{ 2 } \times \frac{ 1 }{ 2 } \times \frac{ 1 }{ 2 } \times \frac{ 1 }{ 2 } \times \frac{ 1 }{ 2 }\] = ......

OpenStudy (anonymous):

p(heads)=5/2 basically.

OpenStudy (anonymous):

yeah and then you would get there lol ^

OpenStudy (anonymous):

@Ottomaticc are you fresh on your statistics?! I need some help..

OpenStudy (anonymous):

(5,2)/32= 5!/2!=60/32=.1.875??

OpenStudy (anonymous):

@hmdp4 My statistics aren't all that fine tuned, but I will be glad to at least look at it and help if I can.

OpenStudy (anonymous):

@Ottomaticc did you get my post?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

i'm going to bump it to the top of the feed right now.